OpenStudy (anonymous):

Solve 7x – 2y = –3 14x + y = 14

OpenStudy (anonymous):

If you put it into a graphing utility you find that it's about (0.75, 4)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

Well, times the top equation by negative two so you get " -14x + 4y = 6" Add both equations together by adding each row of X's, Y's and numbers. -14x + 4y = 6 14x + y = 14 that leaves you with 5y = 20 so y =4 Now just sub y back in to one of the equations to find x :)